### v1

# Hybrid Architecture for Real-Time Video Anomaly Detection: Integrating Spatial and Temporal Analysis

Hybrid Architecture for Real-Time Video Anomaly Detection: Integrating Spatial and Temporal Analysis
Fabien Poirier Paris 8 University, LIASD, France
1 Abstract
We propose a new architecture for real-time anomaly detection in video data, inspired by human behavior by combining spatial and temporal analyses. This approach uses two distinct models: for temporal analysis, a recurrent convolutional network (CNN + RNN) is employed, associating VGG19 and a GRU to process video sequences. Regarding spatial analysis, it is performed using YOLOv7 to analyze individual images. These two analyses can be carried out either in parallel, with a final prediction that combines the results of both analyses, or in series, where the spatial analysis enriches the data before the temporal analysis. In this article, we will compare these two architectural configurations with each other, to evaluate the effectiveness of our hybrid approach in video anomaly detection.
Keywords: Anomaly Detection, Real-Time, YOLOv7, VGG-GRU, Video Analysis, Modular Architecture
2 Introduction
Anomaly detection in videos is essential for various applications, ranging from security surveillance to disaster management, and the monitoring of large-scale events such as the Olympic Games. In such situations, the ability to quickly identify anomalies can have significant consequences, whether it is ensuring the safety of people or responding effectively to critical situations. However, traditional detection systems often rely on an isolated analysis of the temporal aspects of videos, limiting their ability to detect anomalies efficiently in complex environments.
Since videos are multimodal, incorporating both static visual information (images) and dynamic information (sequences of images), it is necessary to adopt an approach capable of capturing the full richness of potential anomalies as efficiently and quickly as possible. To this end, we propose a hybrid architecture combining spatial analysis, for detecting objects and visual patterns in each image, with temporal analysis to model the dynamics of video sequences. For this purpose, we utilize YOLOv7, a state-of-the-art object detection model, coupled with a combination of VGG19 and GRU to model temporal sequences. This approach enables not only the detection of anomalies based on the presence of suspicious objects but also the identification of abnormal behaviors over time.
The originality of our approach lies in the integration of these two types of analysis. We explore two main configurations: a parallel approach (where spatial and temporal analyses are performed simultaneously and combined for a final prediction) and a sequential approach (where spatial analysis enriches the temporal analysis). This integration enhances the precision and reliability of anomaly detection by leveraging the strengths of each type of analysis.
This paper aims to evaluate the effectiveness of these hybrid configurations by comparing the parallel and sequential approaches, while also analyzing the impact of spatial analysis integration into our architecture. We trained and tested both approaches on a proprietary dataset. The results allow us to assess the specific contribution of spatial analysis, highlighting its strengths and limitations within the context of anomaly detection.
Related Work
Anomaly detection in videos has evolved considerably in recent years, thanks to improvements in deep learning technologies and data analysis techniques. This state of the art focuses on recent approaches in object and anomaly detection, as well as hybrid models that combine spatial and temporal analyses.
2.1 Object Detection
The evolution of object detection models has played a crucial role in improving anomaly detection systems. Zou et al. (2019) published a detailed review of the evolution of object detection techniques over the past two decades [25]. Notable advancements include the Recurrent Convolutional Neural Network (RCNN) [8], fast RCNN [7], and faster RCNN [16], as well as the You Only Look Once (YOLO) technology [14, 15, 13]. Successive versions of YOLO have brought significant improvements, with YOLOv4 introducing advanced data augmentation techniques [3], and YOLOv7, which uses the YOLOR (You Only Learn One Representation) architecture [24] for ultra-fast image analysis [23].
2.2 Anomaly Detection
Anomaly detection technologies have also diversified. Doshi and Yilmaz (2020) proposed an approach combining YOLOv3 for object detection and Flownet 2 for optical flow feature extraction, with the information then processed by a KNN [5]. Subsequently, they proposed MONAD (Multi-Objective Neural Anomaly Detector), an architecture composed of two main modules. The first is a feature extraction module based on deep learning, which uses a generative adversarial network (GAN) to predict future video frames and calculates the prediction error (MSE). This module also employs a lightweight object detector, YOLOv3, to extract localization information (center and area of the bounding box) and appearance information (class probabilities) of objects detected in each frame. For each object, a feature vector is constructed, combining the prediction error, localization information, and class probabilities. The second module, dedicated to anomaly detection, uses a non-parametric sequential algorithm to analyze feature vectors in real-time. It compares new observations to normal training data using a k-nearest neighbors (kNN) approach [6]. More recently, Mostafa (2022) introduced the AVAD (Autoencoder-based Video Anomaly Detection) method, which uses a convolutional autoencoder to detect abnormal frames and YOLOv5 to identify objects responsible for anomalies [1]. Anomaly detection models based on recurrent neural networks (RNN) such as LSTM and GRU are also widely used. These models have shown their effectiveness in various applications, such as action recognition [17, 18, 2], anomaly detection [10], and violent action detection [21, 22, 4]. 3D convolution, developed by Ji et al. (2012), is another promising technology, offering good results for action recognition [9, 20, 19] and anomaly detection [11].
2.3 Synthesis
Current approaches often combine advanced object detection techniques with temporal analyses to improve the accuracy of anomaly detection. Hybrid models, which integrate spatial and temporal analyses, promise increased performance by capturing both object characteristics and temporal dynamics of videos.
3 Approach
In this article, we propose an architecture composed of two complementary analyses: a spatial analysis and a temporal analysis. Figure 1 illustrates this dual approach and its overall structure. For the spatial analysis, we have opted for the use of YOLOv7, chosen for its remarkable performance in terms of execution speed and detection accuracy. The temporal analysis, on the other hand, relies on a neural network combining VGG19 (Visual Geometry Group) and GRU (Gated Recurrent Unit), also including an MLP (Multi-Layer Perceptron). This model is designed to process sequences of 15 images, each with a size of 112x112 pixels. VGG-GRU has already been used to reach high performances in the detection of anomaly in videos [12]. Figure 2 presents in detail the structure of this temporal analysis model.
These two models have been trained on proprietary datasets specifically designed for our application. YOLOv7 was trained on a dataset comprising 10,000 images of firearms, 2000 images of fires containing flames or smoke, images of people, as well as images unrelated to these objects, to ensure robust and versatile detection. The VGG-GRU model, on the other hand, was trained on a corpus of videos representing three classes of anomalies: fights (978 videos), gunshots (311 videos), and fires (298 videos). These videos were carefully edited to retain only the segments containing anomalies, and videos containing no anomalies were added to represent the“normal“ class. One of the main advantages of this architecture lies in the flexibility of positioning the spatial analysis model, which can be loaded in parallel or in series according to the specific needs of the application.
Figure 1: Proposed Architecture
Figure 2: Structure of the Temporal Analysis Component VGG-GRU.
The GRU layer has a dropout rate set to 50%, as do all Dense layers, which also have L2 regularization fixed at 0.01.
4 Experiments
In this section, we will present various experiments aimed at comparing the performance of our different architectures. We will begin by evaluating the architecture in which our components are arranged in parallel.
4.1 Parallel Architecture
This architecture uses YOLOv7 to detect various key objects related to our anomalies, based on the assumption that the absence of these objects makes the occurrence of the anomaly highly unlikely. We focus on detecting people for the “fight“ anomaly, the presence of flames and smoke for “fire,“ and the IoU between a person and a firearm for the “gunshot“ anomaly. Once predictions are made by our two analysis modules, they are combined using the following logical rule:
  1. 1.
    If our model detects an anomaly, the predicted class will correspond to this anomaly.
  2. 2.
    Otherwise, if a key object is detected, we will predict the anomaly associated with that object. However, in the specific case of a weapon, we will first check if the Intersection over Union (IoU) between the weapon and a person is greater than 0 before predicting the ”gunshot” anomaly.
For the “gunshot“ class, given that it involves two distinct objects and that a firearm only represents a risk when it is within reach of a person, we use the IoU (Intersection over Union) between these two objects when combining the results of our detections. This approach allows for a more precise evaluation of the spatial context of the detected objects. The results of our model are presented in Table 1.
Table 1: VGG-GRU + YOLO Performance
Accuracy Precision Recall F1-Score
78.42% 85.60% 78.42% 81.16%
Confusion matrix (in percent)
Truth Predicted Fight Gunshot Fire Normal
Fight 63.66% 6.58% 1.93% 27.83%
Gunshot 9.94% 66.06% 9.33% 14.67%
Fire 13.66% 15.73% 57.71% 12.9%
Normal 7.43% 5.96% 3.98% 82.63%
4.1.1 Serial Architecture
As part of our serial analysis, we leveraged YOLOV7 to enrich our input data by applying various preprocessing techniques. Our main approach consisted of removing the background from images, leaving only the key objects detected by YOLO. This method is illustrated in Figure 3. The objective of this technique is to focus our model’s attention on relevant elements, thus avoiding it from concentrating on parasitic background movements. During this experimentation, we explored two options for cases where no key objects were detected by YOLO:
  • Retaining the original image without background: The results of this approach are presented in Table 3.
  • Introducing an entirely black image: This alternative is presented in Table 2.
After a comparative analysis of these two approaches, we observed that using a black background seemed to improve the detection of the “normal“ class. However, this improvement came at the expense of accuracy for other classes (fight, fire). This observation led us to formulate the following hypothesis: in certain scenarios, key objects such as flames/smoke or people were present but were not detected by YOLO, or their detections did not reach the set reliability threshold. Consequently, the use of a black background resulted in the loss of this potentially crucial information. For all these experiments, we set YOLO’s confidence rate to 55% and the threshold to 25%.
Figure 3: Example of Mask Generated with YOLO
Given the behavioral nature of certain anomalies, particularly those involving human interactions, we leveraged the flexibility of our architecture to integrate YOLOv7-pose as a replacement for standard YOLOv7. This adaptation allows us to trace the skeleton of each person present on the screen, thus offering a more refined analysis of movements and postures. We experimented with two preprocessing approaches using YOLOv7-pose:
  • Preservation of the background with superimposition of detected skeletons:
    Visual illustration: Figure 5
    Detailed results: Table 5
  • Removal of the background, presenting only the skeletons on a black background:
    Visual illustration: Figure 5
    Detailed results: Table 4
Table 2: Performance for Mask with Black Background
Accuracy Precision Recall F1-Score
75.58% 79.34% 75.58% 76.50%
Confusion matrix (in percent) for mask with black background
Truth Predicted Fight Gunshot Fire Normal
Fight 55.52% 4.43% 1.76% 38.29%
Gunshot 16.20% 41.44% 1.53% 40.83%
Fire 14.09% 12.49% 20.59% 52.83%
Normal 10.55% 2.98% 2.09% 84.38%
Table 3: Performance for Mask without Black Background
Accuracy Precision Recall F1-Score
72.22% 82.37% 72.22% 75.68%
Confusion matrix (in percent) for mask without black background
Truth Predicted Fight Gunshot Fire Normal
Fight 63.06% 9.33% 1.96% 25.65%
Gunshot 25.99% 41.45% 2.29% 30.27%
Fire 19.23% 15.98% 32.44% 32.35%
Normal 15.33% 4.41% 2.07% 78.19%
The results of these two experiments revealed interesting trade-offs. The analysis of poses on a black background significantly improved the detection of anomalies related to human behaviors, such as fights and gunshots, compared to the previous preprocessing. This improvement is explained by the increased focus of the model on people’s movements and interactions, without distraction from the background. However, this approach also led to a notable decrease in the system’s ability to detect fires. The removal of the background effectively eliminated crucial visual information for identifying flames and smoke, rendering these anomalies undetectable. This experiment highlights the importance of a judicious balance between focusing on human behaviors and preserving contextual information from the environment. It also underscores the need for an adaptive approach in anomaly detection, capable of adjusting based on the specific nature of the anomalies to be detected.
Table 4: YOLOv7-Pose + VGG-GRU without Background (3 Classes)
Accuracy Precision Recall F1-Score
87.2% 90.9% 87.8% 89%
Confusion Matrix (in Percentage) for Pose Analysis without background
Truth Predicted Fight Gunshot Normal
Fight 64.5% 1% 34.5%
Gunshot 11.7% 70.7% 17.6%
Normal 9.5% 0% 90.5%
Table 5: YOLOv7-Pose + VGG-GRU with Background (4 Classes)
Accuracy Precision Recall F1-Score
87.3% 87.6% 87.3% 87.1%
Confusion Matrix (in Percentage) for Pose Analysis without Black Background
Truth Predicted Fight Gunshot Fire Normal
Fight 60.5% 2.4% 1.3% 35.8%
Gunshot 10% 55.6% 14.8% 19.6%
Fire 15.5% 10.6% 48% 25.9%
Normal 3.4% 0.6% 1% 95%
Figure 4: Pose Estimation by YOLOv7 with background
Figure 5: Pose Estimation by YOLOv7 without background
Given the potential impact of the anomalies discussed in this article, we aimed to conclude our experiments by measuring various processing times. Each of these measurements was conducted on a laptop equipped with 32 GB of RAM, an Intel Core i9 processor with 16 cores clocked at 2.3 GHz, and an Nvidia GeForce RTX2080 GPU with 8 GB of dedicated memory. As expected, the parallel architecture proved to be significantly faster than the sequential one (tables 6, 7). This can be explained by the fact that, in the sequential setup, YOLOv7 must pre-process each video frame before being analyzed by our VGG-GRU model. However, the average prediction time for the sequential architecture remains reasonably fast, with an average of 1 to 1.5 seconds per prediction, allowing for a timely response when an anomaly is detected.
It is worth noting that the execution speed of the parallel model can be further improved by adjusting the number of frames YOLOv7 needs to analyze. Due to the combination of both models, it is not necessary to analyze every frame in a sequence to make a prediction. Because of the redundant information between successive frames, the model can be configured to analyze a reduced number of frames per sequence, which enhances processing speed.
Table 6: Execution time of YOLO and CGRU in parallel
Video duration Average detections Processing time
16s 601ms 15s
44s 533ms 35s
9s 994ms 12s
35s 1.1s 57s
23s 1s06 35s
1min 43 758ms 116s (1min 56)
50s 826ms 61s
1min 30 886ms 83s (1min 23)
2s 847ms 847ms
9s 870ms 11s
2s 1s 1s
Table 7: Execution time of YOLO and CGRU in serie
Video duration Average detections Processing time
16s 1s 26s
44s 1s 71s (1min 11)
9s 1.5s 20s
35s 1.5s 81s (1min 21)
23s 1.5s 48s
1min 43 1.2s 193s (3min 13)
50s 1.3s 102s (1min 42)
1min 05 1.4s 134s (2min 14)
2s 1.3s 1.3s
9s 1.3s 17s
2s 1.5s 1.5s
5 Conclusion
Our various experiments have allowed us to identify two optimal configurations for video anomaly detection, each addressing specific requirements:
  • For precise anomaly detection:   The serial configuration, combining YOLOv7 with VGG-GRU, has proven particularly effective. This approach excels in identifying human behavioral anomalies. The integration of pose estimation preprocessing and background removal has significantly improved results, offering detailed analysis adapted to situations where accuracy is paramount.
  • For real-time analysis:
    The parallel architecture, combining YOLOv7 and our VGG-GRU, offers an optimal balance between reliability and speed. This configuration ensures prompt detection of anomalies while maintaining adequate accuracy, thus meeting the needs of applications requiring instant processing.
These results highlight the importance of an adaptive approach in video anomaly detection. Our modular architecture effectively responds to various scenarios, paving the way for diverse applications in the fields of security, surveillance, and event management. The flexibility of our system allows for prioritizing either precision or speed, depending on the specific requirements of each application.
5.1 Limitations and Future Work
Our study has highlighted several challenges and improvement perspectives for our anomaly detection system. Firstly, we observed that not all anomalies are necessarily linked to identifiable key objects, particularly in the case of natural disasters. Moreover, the presence of key objects does not always signify danger, as illustrated by the example of armed military personnel in airports. To overcome these limitations, we are considering directly transmitting the information collected by YOLO and VGG+GRU to our MLP, allowing it to automatically learn the conditions for anomaly detection. We also noticed confusion between certain anomaly classes, suggesting the potential benefit of exploring a binary model to verify this hypothesis. Another promising avenue would be to combine several different processing methods, particularly by only showing detected objects and representing people by their skeletons, which could improve the reliability of the serial model, albeit at the expense of detection speed. Finally, to enhance the robustness and versatility of our system, it would be wise to enrich our dataset with new anomaly classes and compare our architecture to other existing models in the field.